x =2y +7 3x - 2y = 3
Answer by TimothyLamb(4379)

x = 2y + 7
3x - 2y = 3
---
put the system of linear equations into standard form
---
x - 2y = 7
3x - 2y = 3
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -2
y = -4.5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---